OpenStudy (anonymous):

which of the folowing numbers are prime : 2, 11, 1, 6, 21, 9

OpenStudy (cggurumanjunath):

A COMPOSITE NUMBER HAS MORE THAN 2 FACTORS.

OpenStudy (cggurumanjunath):

6=2*3 21=7*3 9=3*3

OpenStudy (cggurumanjunath):

6,21,9 ARE COMPOSITE NUMBERS.

OpenStudy (cggurumanjunath):

2=2*1 11=11*1 2 AND 11 ARE PRIME.

OpenStudy (cggurumanjunath):

1 IS NEITHER PRIME NOR COMPOSITE.

OpenStudy (anonymous):

2 and 11 are the only prime numbers because 1 is not a prime number and all of the other numbers can be divided by something other than one and themselves.

OpenStudy (anonymous):

Alright thank you guys for the help.

OpenStudy (cggurumanjunath):

:)